OpenStudy (anonymous):

For what values of the variable x is the rational expression undefined: 3/(x - 5)

OpenStudy (zpupster):

if a denominator = 0 it is undefined

OpenStudy (cwrw238):

for what value of x is the denominator = 0?

OpenStudy (anonymous):

i think so

OpenStudy (cwrw238):

x - 5 = 0 so what is value of x?

OpenStudy (anonymous):

0

OpenStudy (cwrw238):

???

OpenStudy (cwrw238):

if x -5 =0 then x = 5 as 5-5 = 0

OpenStudy (anonymous):

oh